An aircraft engine operates on a simple ideal Brayton cycle with a pressure ratio of 14.2. Heat is added to the cycle at a rate of450 kW; air passes through the engine at a rate of 1.20 kg/s; and the air at the beginning of the compression is at70 kPa and 0°C. Determine the power produced by this engine and its thermal efficiency. Use constant specific heats at room temperature. Wnet kW Round your final answer to one decimal place. Round your final answer to one decimal place.
Answer #1

P005 K K 1.4 (Tablu A-2a) (213x) (14) 212 斗 =582.6K 千50 1,20 315 Cg 532.6 十 1-005 955.73 K The bmpat th, exit 익 the turbine 터샤 = 1.005 (955.73-447-g) 510,5 KJ He () -1.005 ( 582.6-213) 14-8 81), iw = . 2.0 ( 510.5-311-15) 23 912kW 2311. 32 in450 .53